The coordinate of a particle in meters is given by x(t) = 16t - 3.0t^3, where the time t is in seconds. The particle is momentarily at rest at t=?
1
Expert's answer
2019-02-17T09:16:59-0500

The particle is momentarily at rest when its speed is 0. How to find the speed? Derive it by taking a derivative of coordinate:


"v(t)=x'(t)=(16t - 3.0t^3)'=16-3\cdot 3.0t^{3-1}=0,"


hence:


"0=16-9t^2,""t=\sqrt{16/9}=4/3\text{ s}."